OpenStudy (lifeisadangerousgame):

what's log base 5 of 125?

OpenStudy (anonymous):

thats mean : 5^x=125 ,5^x=5^3 therefore x=3 need more explanation ?

OpenStudy (lifeisadangerousgame):

Nope! That's all, I just asked that cuz my friend who's in a higher grade than me asked me it and wanted to prove I couldn't answer it >:D

OpenStudy (anonymous):

lol,now go and answer it xD ;)

OpenStudy (lifeisadangerousgame):

Mwuahahaha! Thanks!

OpenStudy (anonymous):

xD ,yourwelcome anytimeeee xD